OpenStudy (anonymous):

Write a sequence that has two geometric means between 48 and -750

OpenStudy (anonymous):

\[\frac{x}{48}=\frac{-750}{x}\]

OpenStudy (anonymous):

now cross and time czn u do from here?

OpenStudy (anonymous):

I'm still confused.

OpenStudy (anonymous):

I be back

OpenStudy (anonymous):

\[x^2=-36000\]\[x=\sqrt{-36000}\]\[x=60\sqrt{10i}\]\[or:x= 60i \sqrt{10}\]